QUESTION THREE

Recently, the Transport Association of Zambia increased the price of bus fares for all long routes in Zambia. Specifically, the price of tickets between Lusaka and Livingstone increased from 200 Kwacha to 250 kwacha. The demand equation  is representative of the demand for bus tickets between Lusaka and Livingstone.

a)Calculate the old price quantity demanded level 

b)Calculate the new price quantity demanded level                                                 

c)Work out the Price Elasticity of Demand for tickets from Lusaka to Livingstone between 200 kwacha and 250 kwacha

d)Briefly discuss the Price Elasticity of Demand for tickets from Lusaka to Livingstone between 200 kwacha and 250 kwacha and clarify revenue

   



1
Expert's answer
2021-09-06T13:26:07-0400

Total problem


"Q_d=2000-5P"

a) When P=200 kwacha, "Q_d=2000-5 \times 200=1000"

Old price quantity demanded=1000

b) When P=250 kwacha, "Q_d=2000-5 \times 250=750"

New price quantity demanded=1500

c) Price elasticity of demand "=\frac{\frac{(Q_2-Q_1)}{(Q_1+Q_2)/2}}{\frac{(P_2-P_1)}{(P_2+P_1)/2}}"

"= \frac{\frac{(750-1000)}{(750+1000)/2}}{\frac{(250-200)}{(250+200)/2} } \\

=-1.3"

d) Price Elasticity of Demand for tickets from Lusaka to Livingstone between 200 kwacha and 250 kwacha is elastic. Since the absolute value of elasticity is more than 1. (1.3>1).

So rise in price will lead to fall in total revenue. As rise in price will lead to more fall in quantity demanded. So total revenue will fall.